OpenStudy (anonymous):

What is the first step in solving 3x - 19 = 28? A. Divide by 3 on both sides. B. Multiply by 3 on both sides. C. Add 19 on both sides. D. Subtract 28 on both sides.

OpenStudy (anonymous):

itwas C

OpenStudy (anonymous):

It was C 3x - 19 = 28 19 -19 ------------- 0 9 3x = 9 9/3x x = 3